### v1

# STAPO: Stabilizing Reinforcement Learning for LLMs by Silencing Rare Spurious Tokens

Abstract
Reinforcement Learning (RL) has significantly improved large language model reasoning, but existing RL fine-tuning methods rely heavily on heuristic techniques such as entropy regularization and reweighting to maintain stability. In practice, they often experience late-stage performance collapse, leading to degraded reasoning quality and unstable training. We derive that the magnitude of token-wise policy gradients in RL is negatively correlated with token probability and local policy entropy. Building on this result, we prove that training instability is driven by a tiny fraction of tokens, approximately 0.01%, which we term spurious tokens. When such tokens appear in correct responses, they contribute little to the reasoning outcome but inherit the full sequence-level reward, leading to abnormally amplified gradient updates. Motivated by this observation, we propose Spurious-Token-Aware Policy Optimization (STAPO) for large-scale model refining, which selectively masks such updates and renormalizes the loss over valid tokens. Across six mathematical reasoning benchmarks using Qwen 1.7B, 8B, and 14B base models, STAPO consistently demonstrates superior entropy stability and achieves an average performance improvement of 7.13% over GRPO, 20-Entropy and JustRL.
[…]
Figure 1: (a) Conceptual analogy: We argue that spurious tokens, which are rare and uninformative tokens within otherwise correct responses that receive disproportionately large gradient updates, can harm training stability, analogous to a dissonant vocalist disrupting the harmony of a performance. (b) By masking this negligible fraction (near $0.01\%$) of spurious tokens during the RL process of Qwen3-8B-Base, STAPO approaches the Pareto frontier of performance (AIME24 Acc) and entropy stability, compared to GRPO [1], 20-Entropy [2], and JustRL [3].
1 Introduction
Recent large language models (LLMs), including OpenAI-o1 [4], DeepSeek-R1 [5], and Qwen3 [6], have demonstrated remarkable capabilities in complex reasoning domains such as mathematics and programming. Central to this success is Reinforcement Learning (RL), which optimizes directly for outcome-level correctness and has been empirically linked to the emergence of advanced, long-horizon reasoning behaviors [7]. However, training stability remains a critical bottleneck in RL. Models often suffer from catastrophic degradation, abruptly collapsing from coherent reasoning chains into shallow, repetitive, or nonsensical patterns [8].
Prior work has approached this instability from two perspectives: entropy regulation and gradient modulation. On the entropy front, methods such as selective regularization [2], sample augmentation [9], and clipping modifications [10, 11, 12] attempt to arrest entropy collapse. However, these interventions often precipitate the opposite mode, inducing oscillatory or excessive entropy that degrades reasoning coherence [13]. Turning to the gradient modulation perspective, the disproportionate impact of rare tokens on parameter updates has motivated efforts to suppress unstable gradients via advantage reweighting [14], probability shaping [15], or spark-preserving regularization [16]. Nevertheless, existing approaches lack a fine-grained diagnosis of token instability. By overlooking the interplay between token probability and policy entropy, they fail to distinguish varying degrees of model confidence within low-probability regions, conflating valuable exploration with detrimental noise and leaving the structural sources of instability.
In this work, we trace RL instability to a token-level optimization mechanism. Through a systematic analysis of token updates along the dimensions of entropy, probability, and advantage sign, we uncover a pathological update regime driven by a small subset of spurious tokens. As illustrated in Figure 1, these tokens are rare within otherwise correct responses, yet they provide little to no contribution to the reasoning process and may even introduce incorrect or misleading signals. By masking a negligible fraction (about $0.01\%$) of such spurious tokens during training, we substantially stabilize policy entropy and consistently improve performance across benchmarks. Overall, our contributions are summarized as follows:
  • We derive that the magnitude of token-wise policy gradients in RL is negatively correlated with token probability and local policy entropy. We further establish a unified framework for analyzing training stability in the joint space of policy entropy, token probability, and advantage sign.
  • We identify a pathological class of low-probability, low-entropy tokens with positive advantage, termed spurious tokens, whose amplified gradients destabilize optimization. To address this, we propose Spurious-Token-Aware Policy Optimization (STAPO), which masks these tokens (approximately $0.01\%$) and renormalizes the loss over valid tokens.
  • Across six mathematical reasoning benchmarks and three model scales, STAPO substantially stabilizes policy entropy and consistently improves reasoning performance under diverse evaluation settings, demonstrating its robustness and effectiveness in stabilizing RL training.
2 Preliminary
2.1 Problem Formulation
We focus on fine-tuning LLMs using RL for reasoning tasks. Let $\mathcal{D}$ denote a distribution over input prompts. Given an input prompt $\bm{x}\sim\mathcal{D}$, the LLM parameterized by $\theta$ acts as a stochastic policy $\pi_{\theta}$ that autoregressively generates an output sequence $\bm{y}=(y_{1},\dots,y_{T})$. Specifically, at each step $t$, a token $y_{t}\in\mathcal{V}$ is sampled according to $\pi_{\theta}(y_{t}\mid\bm{x},\bm{y}_{<t})$. Supervision is provided via a sparse, sequence-level verifiable reward $R(\bm{x},\bm{y})\in\{-1,1\}$, which evaluates the correctness of the completed sequence $\bm{y}$ using an external verifier (e.g., a code compiler or a math answer checker). The optimization objective is to maximize the expected reward:
$\mathcal{J}(\theta)=\mathbb{E}_{\bm{x}\sim\mathcal{D},\,\bm{y}\sim\pi_{\theta}(\cdot\mid\bm{x})}\left[R(\bm{x},\bm{y})\right].$ (2)
2.2 Group Relative Policy Optimization (GRPO)
We briefly review Group Relative Policy Optimization (GRPO) [17], which estimates advantages without an explicit value function. For each prompt $\bm{x}$, GRPO samples a group of $G$ output sequences $\{\bm{y}_{1},\dots,\bm{y}_{G}\}$ from a reference policy $\pi_{\theta_{\mathrm{old}}}$. The optimization objective is defined as the average clipped surrogate loss over the group:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{\bm{x}\sim\mathcal{D},\{\bm{y}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|\bm{x})}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|\bm{y}_{i}|}\sum_{t=1}^{|\bm{y}_{i}|}\min\bigl(\rho_{i,t}(\theta)\hat{A}_{i},\;\text{clip}(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}\bigr)\right],$ (3)
where $\rho_{i,t}(\theta)$ is the importance sampling ratio. Since rewards are assigned at the sequence level, all tokens within the same sequence share a common training signal $\hat{A}_{i}$, which is computed by normalizing the sequence reward across samples in the same group.
2.3 Clip-Higher and Token Normalization
Building upon the GRPO objective, DAPO [10] proposes four additional training tricks, among which token-level normalization and an asymmetric clip-higher mechanism are shown to be particularly effective for stabilizing RL training [18, 11, 12]. JustRL [3] adopts these two components and achieves strong empirical performance. The corresponding objective is written as
$\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{\bm{x}\sim\mathcal{D},\{\bm{y}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}(\cdot|\bm{x})}}\left[\frac{1}{\sum_{i=1}^{G}|\bm{y}_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|\bm{y}_{i}|}\min\left(\rho_{i,t}(\theta)\hat{A}_{i},\text{clip}(\rho_{i,t}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}})\hat{A}_{i}\right)\right],$ (4)
where $\epsilon_{\mathrm{low}}$ and $\epsilon_{\mathrm{high}}$ are the asymmetric clipping parameters introduced by the clip-higher mechanism. Due to its simplicity and strong empirical performance, we adopt this configuration as our baseline objective.
3 Methodology
3.1 Token-Level Optimization Mechanisms in RL Training
Understanding how RL-style optimization updates reasoning models is central to analyzing training stability. Prior work has examined entropy dynamics [19, 12] or gradient behavior separately [14, 16]. In contrast, we provide a unified analysis that jointly characterizes gradient scaling and entropy evolution under group-style updates.
We first recall a fundamental gradient property from [14] that characterizes how token probability modulates gradient magnitude. At decoding step $t$ of sequence $\bm{y}_{i}$, the LLM produces a logit vector $\bm{a}_{i,t}\in\mathbb{R}^{N}$ over the vocabulary $\mathcal{V}$, which induces the policy distribution $\pi_{\theta}(\cdot\mid\bm{x},\bm{y}_{i,<t})$ via softmax. We analyze the per-token gradient associated with $y_{i,t}$ as it propagates to intermediate layers.
Theorem 3.1.
(Policy Gradient Norm Bounds). Consider the optimization objective at step $t$ for sample $i$ with target token $y_{i,t}$. The squared $\ell_{2}$-norm of the gradient $\nabla_{\bm{a}}\mathcal{J}$ w.r.t. the logits $\bm{a}$ is directly bounded by the entropy $\mathcal{H}(\pi_{\theta})$ and the target probability $\pi_{\theta}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})$ as follows:
$|w_{i,t}|^{2}\left(1-2\pi_{\theta}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})+e^{-\mathcal{H}(\pi_{\theta})}\right)\leq\|\nabla_{\bm{a}}\mathcal{J}(y_{i,t})\|^{2}\leq|w_{i,t}|^{2}\left(2-2\pi_{\theta}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})-C_{V}\mathcal{H}(\pi_{\theta})^{2}\right),$ (5)
where $C_{V}=\frac{|\mathcal{V}|-1}{|\mathcal{V}|(\ln|\mathcal{V}|)^{2}}$ and the weight $w_{i,t}$ is defined as:
$w_{i,t}=\begin{cases}0,&\text{if }(\hat{A}_{i}>0\land r_{i,t}>1+\epsilon)\lor(\hat{A}_{i}<0\land r_{i,t}<1-\epsilon),\\ \frac{\pi_{\theta}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})}\hat{A}_{i},&\text{otherwise}.\end{cases}$ (6)
Proof.
See Appendix 8 for the detailed derivation. ∎
Theorem 3.1 indicates that the gradient lower bound scales inversely with the token probability $\pi_{\theta}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})$ and the policy entropy $\mathcal{H}(\pi_{\theta})$. Therefore, a token characterized by both low probability and low entropy will lead to a larger gradient norm, which is verified by Figure 2(b). Beyond gradient magnitude, the behavior of LLMs is closely linked to the evolution of policy entropy, which characterizes uncertainty over action distributions. Recent analyses establish a direct connection between policy updates and entropy dynamics.
Lemma 3.2 (Entropy Update Mechanism [19]).
Consider the language policy $\pi_{\theta}(\cdot)$ updated via a natural policy gradient step, the entropy $\mathcal{H}(\pi_{\theta})$ between two consecutive policy iterates satisfies
$\mathcal{H}(\pi_{\theta_{k+1}}(\cdot\mid\bm{x},\bm{y}_{i,<t}))-\mathcal{H}(\pi_{\theta_{k}}(\cdot\mid\bm{x},\bm{y}_{i,<t}))\approx-\eta\,\mathrm{Cov}_{y_{i,t}\sim\pi_{\theta_{k}}(\cdot\mid\bm{x},\bm{y}_{i,<t})}\!\left(\log\pi_{\theta_{k}}(\cdot\mid\bm{x},\bm{y}_{i,<t}),\,\hat{A}_{i}\right).$ (7)
The change in policy entropy can be approximated by the interaction between token probability and advantage, as characterized by Lemma 3.2. Several prior works [12, 20, 13] adopt this perspective to explain entropy dynamics during RL training. However, these approaches mainly regulate entropy through sample selection or reweighting, without fully analyzing the underlying causes of token-level instability. Complementary to this dynamical view, prior studies have observed that learning efficacy is not uniform across the state space, but is concentrated in regions of high policy uncertainty.
Lemma 3.3 (Entropy-Conditioned Learning Potential [2]).
In RL, low-entropy tokens $\mathcal{H}(\pi_{\theta_{k+1}}(\cdot\mid\bm{x},\bm{y}_{i,<t}))<\tau_{h}$ correspond to regions where the policy is already confident and well-optimized, so further updates yield limited learning benefit. In contrast, high-entropy tokens $\mathcal{H}(\pi_{\theta_{k+1}}(\cdot\mid\bm{x},\bm{y}_{i,<t}))\geq\tau_{h}$ mark decision-uncertain regions where multiple plausible continuations exist, and thus constitute the primary loci where policy learning remains effective.
Proposition 3.1 and Lemmas 3.23.3 jointly characterize gradient norm, entropy change, and learning potential. Based on these results, we summarize the optimization impact of token-level updates using the following criteria.
Criteria 1 (Optimization Impact Criteria).
We evaluate the impact of token-level updates :
  1. 1.
    Gradient Norm: Amplified gradients can dominate updates and lead to optimization instability.
  2. 2.
    Entropy Change: Training in RL typically drives entropy upward; updates that further increase entropy can accelerate entropy explosion and distributional drift.
  3. 3.
    Learning Potential: Updates in low-entropy tokens yield limited benefit, as these are already well pretrained.
Table 1: Token Update Phase Diagram. Left: token properties. Right: influence on optimization. Green indicates beneficial updates, and red indicates detrimental updates.
Token Properties Influence on Optimization
Token Prob. Advantage Entropy Gradient Norm Entropy Change [19] Learning Potential [2]
$\pi_{t}\geq\tau_{p}$ $\hat{A}<0$ $\mathcal{H}\geq\tau_{h}$ Low $\uparrow$ High
$\pi_{t}\geq\tau_{p}$ $\hat{A}>0$ $\mathcal{H}\geq\tau_{h}$ Low $\downarrow$ High
$\pi_{t}<\tau_{p}$ $\hat{A}<0$ $\mathcal{H}\geq\tau_{h}$ High $\downarrow$ High
$\pi_{t}<\tau_{p}$ $\hat{A}>0$ $\mathcal{H}\geq\tau_{h}$ High $\uparrow$ High
$\pi_{t}\geq\tau_{p}$ $\hat{A}<0$ $\mathcal{H}<\tau_{h}$ Low $\uparrow$ Low
$\pi_{t}\geq\tau_{p}$ $\hat{A}>0$ $\mathcal{H}<\tau_{h}$ Low $\downarrow$ Low
$\pi_{t}<\tau_{p}$ $\hat{A}<0$ $\mathcal{H}<\tau_{h}$ High $\downarrow$ Low
[…]
(a) Token Quadrants Visualization
(b) Gradient Norm Comparison
(c) Illustrative examples of spurious tokens. Three instances where the LLM selects spurious tokens (pink) over reasonable top-2 candidates (green). Spurious tokens often represent semantic errors despite the final response correctness. More comprehensive examples are provided in Appendix 9.
Figure 2: Comprehensive Analysis of Spurious Tokens.
3.2 Spurious Tokens as a Destructive Update Mode
To systematically analyze how different token types influence optimization, we categorize tokens along three binary factors: state entropy (high/low), token probability (high/low), and advantage sign (positive/negative). We then interpret the optimization behavior of each regime using Criteria 1, as summarized in Table 1. We observe that tokens characterized by low entropy, low probability, and positive advantage lead to negative optimization effects across all evaluation criteria.
To empirically validate this hypothesis and examine the real-world behavior of such tokens, we fine-tune Qwen3-8B [6] under the JustRL setting [3] and explicitly record the token statistics during training. As visualized in Figure 2(a), the recorded tokens naturally form distinct clusters in the probability-entropy plane, confirming the existence of the four hypothesized regimes. Crucially, statistical analysis of these recorded dynamics (Figure 2(b)) reveals that spurious tokens trigger the highest mean gradient norms among all categories (e.g., $16.7\%$ relative to the high-entropy & high-probability baseline). This quantitative evidence suggests that when these rare, low-entropy errors occur, they result in disproportionately large model updates.
Qualitative examples in Figure 2(c) further illustrate that these tokens are often semantically inappropriate or erroneous breaks in coherence, whereas top candidates represent more reasonable choices. Since the final response may still receive a high reward, the large gradient updates observed in Figure 2(b) significantly reinforce these erroneous patterns. Building on these empirical observations, we formally define this phenomenon under sequence-level reward assignment:
Definition 3.4 (Spurious Tokens).
Spurious tokens are tokens that contribute negligibly to the correctness of a reasoning outcome but receive disproportionately large positive updates due to sequence-level reward assignment.
This over-amplification of semantic errors disrupts chain-of-thought coherence and ultimately degrades training stability. As demonstrated by the alignment between the statistical anomalies (Figure 2(b)) and semantic errors (Figure 2(c)), the condition of low entropy, low probability, and positive advantage serves as a robust practical criterion for identifying spurious tokens.
3.3 Spurious-Token-Aware Policy Optimization (STAPO)
To mitigate the detrimental effects of spurious tokens while preserving the policy’s capacity for legitimate error correction, we propose Spurious-Token-Aware Policy Optimization (STAPO). The core of STAPO is the Silencing Spurious Tokens (S2T) mechanism, which selectively discards policy gradient contributions from spurious tokens. Specifically, S2T introduces a binary mask $\mathbb{I}^{\text{S2T}}_{i,t}$ by suppressing gradient contributions from tokens that fall into the destructive regime:
$\mathbb{I}^{\text{S2T}}_{i,t}=\begin{cases}0,&\text{if }\hat{A}_{i}>0\land\pi(y_{i,t})<\tau_{p}\land\mathcal{H}_{t}<\tau_{h},\\ 1,&\text{otherwise}.\end{cases}$ (8)
Here, $\tau_{p}$ and $\tau_{h}$ are hyperparameters representing the probability and entropy thresholds, respectively. Then the S2T mask is applied to the token-level policy gradient, yielding the STAPO objective:
$\mathcal{J}_{\text{STAPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}\bm{x}\sim\mathcal{D}\\ \{\bm{y}_{i}\}_{i=1}^{G}\sim\pi_{\mathrm{old}}\end{subarray}}\left[\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sum_{i=1}^{G}\sum_{t=1}^{|\bm{y}_{i}|}\mathbb{I}^{\text{S2T}}_{i,t}}}\sum_{i=1}^{G}\sum_{t=1}^{|\bm{y}_{i}|}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{I}^{\text{S2T}}_{i,t}}\cdot\min\left(\rho_{i,t}\hat{A}_{i},\text{clip}(\rho_{i,t},1-\epsilon,1+\epsilon)\hat{A}_{i}\right)\right].$ (9)
The terms $\rho_{i,t}$ and $\hat{A}_{i}$ follow the standard definitions for probability ratio and group-normalized advantage, respectively:
$\rho_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})},\quad\hat{A}_{i}=\frac{R(\bm{x},\bm{y}_{i})-\mathrm{mean}(\{R(\bm{x},\bm{y}_{j})\}_{j=1}^{G})}{\mathrm{std}(\{R(\bm{x},\bm{y}_{j})\}_{j=1}^{G})},$ (10)
Comparing Eq. (9) with the standard DAPO objective, there are two main differences highlighted in red. Firstly, STAPO selectively masks the loss calculation for spurious tokens. Furthermore, the normalization term in Eq. (9) is adjusted to account only for the remaining valid tokens. This design effectively prevents the misleading update due to the stochastic sampling process, thereby stabilizing the RL alignment process. The overall algorithm is summarized in Algorithm 1.
Remark 3.5 (Thresholding Strategy).
Following [2], the entropy threshold $\tau_{h}$ is implemented as a dynamic quantile within each mini-batch to adaptively capture tokens with low epistemic uncertainty. In contrast, we purposefully employ a fixed absolute value for the probability threshold $\tau_{p}$ rather than a quantile. Our empirical findings indicate that applying a quantile-based $\tau_{p}$ would systematically discard a fixed proportion of tokens regardless of their absolute confidence, which inevitably prunes legitimate high-probability tokens and compromises training stability.
Algorithm 1 Spurious-Token-Aware Policy Optimization (STAPO)
1:Dataset $\mathcal{D}$, Initial Policy $\pi_{\theta}$, Group size $G$, Thresholds $\tau_{p},\tau_{h}$, Batch size $B$
2:Initialize policy parameters $\theta$
3:for each training iteration do
4:  Synchronize: $\theta_{\text{old}}\leftarrow\theta$
5:  Sample prompts $\bm{x}^{B}\sim\mathcal{D}$ and generate responses $\{\bm{y}_{1},\dots,\bm{y}_{G}\}^{B}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid\bm{x}^{B})$
6:  Compute advantages $\hat{A}^{B}_{i}$ using Eq. (10)
7:  for each mini-batch $\subset\text{Batch Data}$ of size $B$ do
8:   for each response $\bm{y}_{i}$ and token $t$ do
9:     Obtain $p_{i,t}=\pi_{\theta}(y_{i,t}\mid\bm{x},\bm{y}_{i,<t})$ and $h_{i,t}=\mathcal{H}(\pi_{\theta}(\cdot\mid\bm{x},\bm{y}_{i,<t}))$
10:     $\mathbb{I}^{\text{silence}}_{i,t}\leftarrow 1$ $\triangleright$ Default: Keep
11:     if $\hat{A}_{i}>0\land p_{i,t}<\tau_{p}\land h_{i,t}<\tau_{h}$ then $\mathbb{I}^{\text{S2T}}_{i,t}\leftarrow 0$ $\triangleright$ Spurious Tokens: Mask
12:   end for
13:   Update $\theta$ through Eq. (9) with $\mathbb{I}^{\text{S2T}}_{i,t}$
14:  end for
15:end for
16:return Final policy $\pi_{\theta}$
4 Related Work
Reinforcement Learning for LLMs. In contrast to traditional reinforcement learning paradigms applied to control and decision-making problems [21, 22, 23], reinforcement learning from human feedback (RLHF) has emerged as a predominant approach for aligning LLMs with human preferences and diverse downstream objectives [4, 5]. While early approaches relied on on-policy optimization such as PPO [24], recent work has shifted toward more efficient preference-based DPO [25], which avoids explicit reward modeling and online rollouts. The focus has further expanded from general alignment to improving reasoning capabilities, motivating RL and training schemes such as GRPO [17]. In this context, a range of policy optimization variants, including DAPO [10], GSPO [18], SAPO [11], and other related methods [26, 27, 9], have been developed to enhance optimization stability, sample efficiency, and scalability for reasoning-oriented language models.
Entropy Instability in RL. A persistent challenge in training reasoning-oriented language models with RL is the rapid collapse of policy entropy during early optimization. Prior work mitigates this issue through interventions such as selectively regularizing high-entropy tokens [2], increasing the proportion of entropy-enhancing samples [12], and modifying clipping strategies in policy optimization [10, 28, 11]. However, these methods often introduce the opposite failure mode, where entropy grows excessively or becomes unstable, degrading reasoning coherence and leading to repetitive or unstructured outputs. Although some studies analyze entropy dynamics during RL training [19, 20] and others explicitly enforce entropy stability as an optimization objective [13], existing approaches largely treat entropy as a surface-level training signal rather than addressing the underlying sources of instability.
Gradient Domination by Low-Probability Tokens. A key microscopic source of instability stems from the disproportionate gradient influence of low-probability tokens. As proved by [14], rare tokens can generate excessively large updates, allowing a small subset of unstable predictions to dominate optimization, which harms fine-tuning stability. Recent work addresses this through probability-aware modulation. Yang et al. [14] propose Advantage Reweighting and Low-Probability Token Isolation to attenuate or asynchronously update rare tokens, while Token-Regulated GRPO (TR-GRPO) [15], using probability shaping to suppress sharp directions induced by low-confidence predictions. To avoid suppressing informative exploratory signals, Low-Probability Regularization (Lp-Reg) [16], which filters noise while preserving meaningful rare tokens, and GTPO [29] mitigates conflicting gradients arising from such tokens across reward trajectories. However, these approaches largely rely on scalar probability thresholds, lacking a joint, fine-grained treatment of token-level confidence and probability, and thus failing to distinguish useful exploration from aleatoric noise under local model calibration.
5 Experiments
5.1 Settings
Baselines. We compare our approach against several state-of-the-art reinforcement learning algorithms for LLMs, including GRPO [17], 20-Entropy [2] and JustRL [3]. For a fair comparison, the dynamic sampling technique introduced in [10] is not applied to any of the baselines.
[…]
Benchmarks. We evaluate our models on six widely adopted and challenging mathematical reasoning benchmarks: AIME24 [32], AIME25 [33], AMC23 [32] , MATH500 [34], Minerva [35], and OlympiadBench [36]. We generate $N$ independent responses per problem ($N=4$ for MATH500, Minerva, and OlympiadBench; $N=32$ for others) across two decoding configurations: temperature=1.0, top-p=1.0 and temperature=0.7, top-p=0.9, with a maximum length of 20,480 tokens. Results are reported as the average accuracy. To ensure evaluation rigour, we employ CompassVerifier-3B [37], a lightweight LLM verifier, to rectify misclassifications from the rule-based verification.
(a) AIME24 Acc, Entropy, and Training Reward for Qwen3-1.7B Base.
(b) AIME24 Acc, Entropy, and Training Reward for Qwen3-8B Base.
(c) AIME24 Acc, Entropy, and Training Reward for Qwen3-14B Base.
Figure 3: Training Results across Different Model Scales. Each row presents the training dynamics for a specific model size (1.7B, 8B, and 14B). Notably, our proposed STAPO method achieves superior performance while maintaining stable policy entropy across all model sizes.
5.2 Main Results
In this section, we present a comprehensive analysis of STAPO across three model scales (1.7B, 8B, and 14B). We structure our analysis into three dimensions: the training dynamics, the performance under training-aligned configurations, and the robustness across different evaluation protocols.
5.2.1 Training Behaviors Analysis
We analyze the training behaviors in Figure 3, where policy entropy serves as a key indicator of RL stability in reasoning tasks. As shown in the middle column, 20-Entropy and JustRL frequently suffer from entropy explosion, while GRPO exhibits entropy collapse, both of which are detrimental to stable training. In contrast, STAPO maintains a stable and well-regulated entropy level across all model scales. Notably, this stability is achieved by masking only a negligible fraction of spurious tokens (approximately 0.01% in our experiments). Moreover, STAPO attains the highest training reward (right column) and steadily improves AIME24 validation accuracy (left column) across all model sizes, suggesting that filtering out spurious tokens mitigates optimization distortions and leads to more stable and reliable learning.
5.2.2 Performance under Training-Aligned Settings
Table 2 summarizes the main results. We first focus on the Training-Aligned evaluation setting (temperature=1.0, top-p=1.0), which reflects the model’s intrinsic capability without the aid of greedy decoding strategies. Under this setting, STAPO demonstrates consistent scalability and superior performance across all model sizes. For the 1.7B model, STAPO achieves a substantial lead, surpassing the strongest baseline (20-Entropy) by a relative margin of 13.50% in average accuracy. At the 8B scale, STAPO continues to outperform all competing methods, maintaining its position as the state-of-the-art. The efficacy of our method is further validated at the 14B scale, where STAPO extends its lead, achieving a relative improvement of approximately 5.94% over the best-performing baseline. These results collectively demonstrate that STAPO effectively unlocks the reasoning potential of base models across varying scales.
5.2.3 Performance Across Evaluation Settings
Recent works adopt different evaluation configurations [10, 2, 3, 12]. To ensure a fair comparison, we additionally evaluate STAPO under the JustRL setting (shown in parentheses/gray in Table 2). The results indicate that STAPO remains the top-performing method across different model scales under this setting. However, the performance gap between STAPO and the baselines becomes slightly narrower compared to the training-aligned configuration. We attribute this effect to the evaluation protocol: baselines with unstable, high-entropy distributions benefit more from constrained decoding, which suppresses low-probability “tail” generations. In contrast, STAPO learns a more intrinsically stable distribution and therefore relies less on decoding heuristics. Overall, STAPO achieves state-of-the-art performance under both configurations, demonstrating its robustness and stability.
Table 2: Main Results on Six Mathematical Reasoning Benchmarks across Three Model Scales. Each cell reports performance under the training-aligned configuration (temperature = 1.0, top-p = 1.0) and the JustRL [3] evaluation setting (temperature = 0.7, top-p = 0.9, shown in gray). For each configuration, the best result is highlighted in bold (training-aligned setting) and in dark gray (JustRL setting), respectively.
Baseline AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
RL from the Qwen3-1.7B Base Model
GRPO 6.88  (10.83) 4.06  (4.06) 40.47  (41.17) 68.60  (67.15) 32.44  (32.53) 30.64  (30.64) 30.52  (31.06)
20-Entropy 13.33  (13.85) 8.44  (7.60) 46.56  (44.53) 69.85  (69.10) 28.03  (29.04) 35.65  (34.35) 33.64  (33.08)
JustRL 8.85  (14.58) 6.25  (10.93) 45.39  (51.09) 62.70  (69.75) 22.61  (28.89) 29.15  (34.16) 29.16  (34.90)
STAPO 17.40  (16.04) 15.42  (14.27) 55.94  (52.42) 73.55  (70.30) 28.22  (29.42) 38.54  (37.87) 38.18  (36.72)
RL from the Qwen3-8B Base Model
GRPO 31.25  (32.40) 24.69  (24.17) 75.23  (74.14) 88.90  (88.85) 55.88  (53.58) 61.50  (58.34) 56.24  (55.25)
20-Entropy 31.25  (35.10) 27.50  (27.39) 79.92  (80.00) 89.85  (89.85) 54.78  (54.23) 62.50  (60.39) 57.63  (57.83)
JustRL 25.21  (34.06) 21.98  (26.04) 73.52  (81.48) 84.90  (87.55) 47.33  (51.03) 51.26  (56.57) 50.70  (56.12)
STAPO 33.44  (39.48) 28.65  (29.37) 79.92  (80.23) 90.40  (91.40) 57.17  (55.70) 62.98  (63.32) 58.76  (59.92)
RL from the Qwen3-14B Base Model
GRPO 42.19  (40.52) 30.31  (32.40) 81.95  (78.12) 90.05  (92.00) 55.15  (57.54) 64.76  (64.28) 60.74  (60.81)
20-Entropy 42.08  (51.25) 34.06  (39.48) 84.45  (87.73) 91.65  (92.30) 51.10  (54.78) 61.30  (63.95) 60.77  (64.92)
JustRL 37.40  (52.08) 26.56  (38.96) 81.95  (89.61) 87.65  (94.25) 43.84  (58.92) 57.01  (68.81) 55.74  (67.11)
STAPO 46.98  (54.27) 35.21  (41.67) 87.11  (90.62) 92.45  (93.85) 59.47  (59.93) 68.21  (71.62) 64.38  (68.66)
(a) Sensitivity to Probability Threshold ($\tau_{p}$)
(b) Sensitivity to Entropy Threshold ($\tau_{h}$)
Figure 4: Sensitivity Analysis of Probability Threshold $\tau_{p}$ and Entropy Threshold $\tau_{h}$ for Qwen3-1.7B Base. Performance is reported as Acc (avg@32) on AIME24 and AIME25 benchmarks.
5.3 Hyperparameter Sensitivity
Sensitivity to Probability Threshold ($\tau_{p}$). We evaluate the sensitivity of the Qwen3-1.7B Base model to the probability threshold $\tau_{p}$. As illustrated in Figure 4(a), accuracy declines sharply as $\tau_{p}$ increases. On AIME24, performance drops from 17.4% to 7.2% as $\tau_{p}$ shifts from $2\times 10^{-3}$ to $2\times 10^{-1}$. This trend indicates that overly aggressive thresholds remove low-probability tokens that are crucial for maintaining valid reasoning chains. Therefore, effective filtering must remain highly selective, targeting only genuinely spurious tokens while preserving rare yet semantically meaningful ones.
Sensitivity to Entropy Threshold ($\tau_{h}$). The threshold $\tau_{h}$ defines the percentile boundary for entropy-based masking. In our optimal configuration ($\tau_{h}=20\%$), we mask the bottom 80% of low-entropy tokens. As $\tau_{h}$ increases, performance consistently declines. Raising $\tau_{h}$ from 20% to 80% reduces AIME24 accuracy from 17.4% to 11.6% and AIME25 accuracy from 15.4% to 7.3%. This degradation suggests that less stringent masking permits more detrimental updates from uninformative tokens, which can bias optimization toward trivial patterns. Overall, these results indicate that a lower $\tau_{h}$ is crucial for stabilizing training.
(a) Average Acc (avg@32) across six benchmarks with temperature=1.0, top-p=1.0.
(b) Average Acc (avg@32) across six benchmarks with temperature=0.7, top-p=0.9.
Figure 5: Ablation Results of Different Masking Strategies across Three Model Scales.
5.4 Masking Strategy Ablation
As illustrated in Figure 5, our ablation study across three model scales reveals distinct performance trajectories contingent upon the masking criteria employed. Specifically, masking tokens based exclusively on a probability threshold consistently underperforms the baseline, suggesting that an indiscriminate pruning of low-probability tokens inadvertently discards infrequent but logically indispensable reasoning steps. In contrast, the strategy of masking high-entropy, low-probability tokens exhibits a significant scale-dependent divergence: while it remains competitive for the 8B and 14B models, it leads to a catastrophic performance collapse on the 1.7B scale. This disparity implies that while larger models may possess the structural redundancy to withstand the removal of high-uncertainty noise, smaller models rely heavily on these tokens for exploratory knowledge representation. Most notably, STAPO is the only approach that achieves a robust performance gain over the baseline across all parameter scales. By precisely isolating these spurious signals, STAPO effectively regularizes gradient updates without compromising the model’s fundamental capacity for legitimate mathematical derivation.
5.5 Detailed Analysis
5.5.1 Spurious Tokens Ratio
As shown in Figure 6(a), the proportion of spurious tokens identified by STAPO remains extremely small throughout RL training. In the early stage, the ratio briefly peaks at approximately 0.03% for the 14B model before rapidly declining. For the majority of training, it consistently stays below 0.01% across all model scales. Despite their rarity, masking these tokens yields substantial performance improvements (see Section 5.2). These results suggest that RL instability is not primarily caused by frequent errors, but rather by a sparse set of positive advantage yet semantically uninformative tokens that exert disproportionate influence on gradient updates.
5.5.2 Word Cloud Visualization
The qualitative distinction between spurious and normal tokens is illustrated in Figures 6(b) and 6(c). The most frequent spurious tokens mainly include specific digits (e.g., “4”, “1”, “2”), mathematical symbols (e.g., “\$”), and transitional words such as “Wait”, “But”, and “Since”. Although these tokens may appear in correct responses, they typically contribute little to the underlying logical derivation. When coupled with large gradient updates, they can induce disproportionately unstable optimization dynamics. In contrast, the retained normal tokens correspond to the structural vocabulary of mathematical reasoning, including words such as “Let”, “find”, “we”, and “can”. These tokens represent essential procedural and logical components that sustain coherent reasoning chains.
(a) Spurious Tokens Ratio.
(b) Frequent Spurious Tokens.
(c) Frequent Normal Tokens.
Figure 6: Analysis of Spurious Tokens: (a) indicates the ratio of spurious tokens during training; (b) and (c) present the word clouds of frequently masked spurious tokens and retained normal tokens
6 Conclusion
In this work, we investigate the instability of RL from a token-level perspective and identify a pathological update regime driven by the interaction between token probability, policy entropy, and advantage sign. We show that a negligible fraction of spurious tokens, defined as rare and uninformative tokens that coincidentally co-occur with correct outcomes, can receive disproportionately amplified gradients and exert an outsized influence on optimization dynamics. Based on this analysis, we propose STAPO, a simple yet effective method that suppresses approxim ately $0.01\%$ of destructive token updates and renormalizes the loss over valid tokens. Empirical results across six mathematical reasoning benchmarks demonstrate that this minimal intervention substantially stabilizes policy entropy and consistently improves reasoning performance.
Due to resource and time constraints, we have not yet conducted ablation studies covering all classification scenarios, and our experiments currently focus primarily on mathematical reasoning tasks, lacking evaluation in other domains such as code generation. Moreover, the current analysis focuses primarily on tokens within correct responses. In future work, we plan to extend both the empirical evaluation to broader task domains and to include tokens from incorrect responses, aiming to develop a more comprehensive understanding of token-level dynamics in RL. We hope this study provides new insights into improving the stability of RL for LLM reasoning.